# Generalized and Unified Equivalences between Hardness and Pseudoentropy

### v1

@@ 2.2 Mirror Descent
[…]
@@
[…]
@@ 2.3 The Multiplicative Weights Algorithm
We give a proof of Lemma 2.3 using Pinsker’s inequality in Section A.1.
@@ 3 Regularity Lemma with Weight-Restricted Calibration
As we discuss in Section 1.5, the key to proving our main theorems (Theorems 1.10 and 1.11) is to establish an enhanced regularity lemma that allows us to construct a low complexity function $s:\{0,1\}^{n}\to\Delta_{L}$ satisfying both multiaccuracy (12) and weight-restricted calibration (11). We formally state and prove this result in both the non-uniform and uniform settings.
@@ 3.1 Non-uniform Setting
We prove Theorem 3.1 by constructing $s$ using Algorithm 1 based on the Multiplicative Weights technique discussed in Section 2.3. In particular, our analysis tracks the Fenchel-Young divergence $\Gamma=\Gamma_{\varphi,\psi}$ defined by the negative Shannon entropy $\varphi$ in (25) and its convex conjugate $\psi$ in (26). For every $g:\{0,1\}^{n}\to\Delta_{L}$ and $h:\{0,1\}^{n}\to\mathbb{R}$, we define
@@
[…]
@@
As we will show, Theorem 3.1 follows from Lemmas 3.2 and 3.3 below.
@@ 18             $\mathtt{updated}\leftarrow\mathtt{TRUE}$;
Algorithm 1 terminates after $k=O((\log L)/\varepsilon^{2})$ updates. Moreover, the output $s$ produced by the algorithm satisfies
@@
Therefore, (28) implies that $k=O((\log L)/\varepsilon^{2})$. Moreover, when Algorithm 1 terminates, both conditions at Algorithms 1 and 1 are violated, implying (27).
It remains to prove (28). We will prove the stronger claim that whenever we update from $h_{k}$ to $h_{k+1}$ at Algorithm 1 or Algorithm 1,
@@
We focus on the case where the update occurs at Algorithm 1, and the other case of Algorithm 1 can be handled similarly. By Lemma 2.2, Lemma 2.3 and the fact that $\|f\|_{\infty}\leq 1$, we have
@@
By the condition at Algorithm 1 and the fact that $\|\hat{g}_{k}-g_{k}\|_{1}\leq\varepsilon/10$, we have
@@
There exists an implementation of Algorithm 1 such that the following holds. Assume that the classes $\mathcal{F}$ and $\mathcal{R}$ both consist of functions with circuit complexity at most $T$. Then for every $k$, the circuit complexity of $\hat{g}_{k}$ is at most $(k+1)(T+L\cdot\mathsf{poly}(1/\varepsilon,\log L))$, and each coordinate of $\hat{g}_{k}(x)$ for every $x\in\{0,1\}^{n}$ is represented in binary format with $O(\log(L/\varepsilon))$ bits.
We will use the following lemma to prove Lemma 3.3:
@@
$v:=\frac{1}{\sum_{y\in[L]}p_{y}}(p_{1},\ldots,p_{L})=\mathsf{softmax}(\tilde{q}).$
Since each $p_{y}$ can be computed in circuit complexity $\mathsf{poly}(B,\log(1/\varepsilon))$, we have that $v$ can be computed in circuit complexity $L\cdot\mathsf{poly}(B,\log(1/\varepsilon),\log L)$. By (32) and the Lipschitzness of $\mathsf{softmax}$ (Lemma 2.3), we have
@@
We are now ready to prove Lemma 3.3.
Proof of Lemma 3.3.
Let $\kappa=O((\log L)/\varepsilon^{2})$ be an upper bound on the total number of updates in Algorithm 1, as guaranteed by Lemma 3.2. Throughout the algorithm, we maintain a circuit $C_{k}$ that computes $x\mapsto(\hat{h}_{k}(x),\hat{g}_{k}(x))$ where $\hat{h}_{k}:\{0,1\}^{n}\to\mathbb{R}^{L},\hat{g}_{k}:\{0,1\}^{n}\to\Delta_{L}$ satisfy the following:
@@
    $\|\hat{g}_{k}-g_{k}\|_{\infty}\leq\varepsilon/10$ as required by Algorithms 1, 1 and 1;
@@
We prove this by induction on $t$. When $t=0$, the algorithm initializes $h_{0}$ to be the constant zero function, so $\hat{h}_{0}$ can be computed in circuit complexity $O(L)$. By Lemma 3.4, $\hat{g}_{0}$ can be computed in circuit complexity $L\cdot\mathsf{poly}(\log(1/\varepsilon),\log L)$.
Now suppose we have a circuit $C_{k}$ with size at most $T_{k}$ that computes $x\mapsto(\hat{h}_{k}(x),\hat{g}_{k}(x))$ satisfying the five conditions above, and let us construct the circuit $C_{k+1}$ that computes $x\mapsto(\hat{h}_{k+1}(x),\hat{g}_{k+1}(x))$. In Algorithm 1, $h_{k+1}$ may be updated from $h_{k}$ at either Algorithm 1 (using $f$) or Algorithm 1 (using $r\circ\hat{g}_{t}$). Both cases can be handled similarly using our assumption that both $f$ and $r$ have circuit complexity at most $T$. We focus on the first case for simplicity. We first compute $\hat{f}$ with $\|\hat{f}\|_{\infty}\leq\varepsilon$ such that $\|\hat{f}-\varepsilon f\|_{\infty}\leq\varepsilon/30\kappa$, where each coordinate of $\hat{f}(x)$ need only have $O(\log(\kappa/\varepsilon))$ bits in binary representation. This can be done in circuit complexity $T+L\cdot\mathsf{poly}(\log(\kappa/\varepsilon))$, noting that we only need to read $O(\log(\kappa/\varepsilon))$ bits from each coordinate of $f(x)$. Then we subtract $\hat{f}$ from $\hat{h}_{k}$ to obtain $\hat{h}_{k+1}$. The subtraction takes circuit complexity $L\cdot\mathsf{poly}(\log(\kappa/\varepsilon))$ and ensures that $h_{k+1}$ satisfies conditions 1,2, and 4. Finally, by Lemma 3.4, we can compute $\hat{g}_{k+1}$ satisfying conditions 3 and 5 from $\hat{f}_{k+1}$ in circuit complexity $O(L\cdot\mathsf{poly}(\kappa\varepsilon,\log(1/\varepsilon),\log L))$. The total circuit complexity of $C_{k+1}$ is
@@
We now complete the proof of Theorem 3.1 using Lemma 3.2 and Lemma 3.3.
Proof of Theorem 3.1.
Lemma 3.2 ensures that the output function $s$ satisfies Conditions 2 and 3 in Theorem 3.1 (indistinguishability and weight-restricted calibration). By Lemma 3.3, we can make sure that $s$ has circuit complexity at most
@@
where we used the fact that $k+1=O((\log L)/\varepsilon^{2})$ by Lemma 3.2. This establishes Condition 1 required by Theorem 3.1. ∎
@@ 3.2 Uniform Setting
In Definitions 3.8 and 3.9 above, the success probability of an oracle can be easily amplified by independent repetition. Suppose we have an oracle with success probability $\alpha>0$. We can construct an oracle with success probability at least $1-\delta$ by independently running the oracle $u=O(\alpha^{-1}\log(1/\delta))$ times to ensure that with probability at least $21-\delta/$ 2, at least one of the runs is successful. We can then perform a simple test using $O(\varepsilon^{-2}\log(u/\delta)))$ additional fresh data points to select a successful run with probability at least $21-\delta/$ 2 (with a constant-factor loss in $\varepsilon$). The overall success probability is at least $1-\delta$ by the union bound.
@@ 26             $\mathtt{updated}\leftarrow\mathtt{TRUE}$;
We prove Theorem 3.10 using Algorithm 2 as the simulating algorithm $\mathcal{S}$ based on the same multiplicative weights idea as Algorithm 1. The difference is that instead of directly looking for $f$ and $r$ as in Lines 1 and 1 of Algorithm 1, we obtain them efficiently by invoking the oracles $\mathcal{A}$ and $\mathcal{B}$ at Lines 2 and 2 in Algorithm 2. Recall that for $y\in[L]$, we use $\mathbf{e}_{y}$ to denote the standard basis vector whose $y$-th coordinate is one. This ensures that for every $v\in\Delta_{L}$, we have $\mathbb{E}_{y\sim v}[\mathbf{e}_{y}]=v$.
Theorem 3.10 follows immediately from the following lemma about Algorithm 2:
@@
For every $\varepsilon\in(0,1/2)$, there exist $\kappa=O((\log L)/\varepsilon^{2}),\delta^{\prime}=\Omega(\delta/\kappa),m^{\prime}=O((\log(1/\delta^{\prime}))/\varepsilon^{2})$ such that when we run Algorithm 2 with parameters $(\varepsilon,m^{\prime})$, with probability at least $1-\delta$, the following holds. Algorithm 2 terminates after $k\leq\kappa$ updates and returns a function $s:\{0,1\}^{n}\to[-1,1]^{L}$ that is simultaneously $(\mathcal{A},\varepsilon,\delta^{\prime})$-indistinguishable from $g^{*}$ and $(\mathcal{B},\varepsilon,\delta^{\prime})$-calibrated w.r.t. $g^{*}$. Also, Algorithm 2 runs in time $\mathsf{poly}(m,T,L,1/\varepsilon,\log(1/\delta))$, and for every $x\in\{0,1\}^{n}$, the function value $s(x)$ can be evaluated in time $\mathsf{poly}(m,T,L,1/\varepsilon,\log(1/\delta))$.
@@
By the Chernoff bound, choosing $m^{\prime}=O((\log(1/\delta^{\prime}))/\varepsilon^{2})$ is sufficient to ensure that each time Algorithm 2 is performed, with probability at least $1-\delta^{\prime}$ we have
@@
Similarly, each time when Algorithm 2 is performed, with probability at least $1-\delta^{\prime}$ we have
@@
By the definition of $(\mathcal{A},\varepsilon,\delta^{\prime})$-distinguishability (Definition 1.6), each time when Algorithm 2 is performed with $\hat{g}_{k}$ being $(\mathcal{A},\varepsilon,\delta^{\prime})$-distinguishable, with probability at least $1-\delta^{\prime}$ we have
@@
Similarly, by the definition of $(\mathcal{B},\varepsilon,\delta^{\prime})$-miscalibration Definition 3.9, each time when Algorithm 2 is performed with $\hat{g}_{k}$ being $(\mathcal{B},\varepsilon,\delta^{\prime})$-miscalibrated, with probability at least $1-\delta^{\prime}$ we have
@@
By the union bound, we can choose $\delta^{\prime}=\Omega(\delta/\kappa)$ to ensure that with probability at least $1-\delta$, all of the above are satisfied simultaneously in the first $\kappa+1$ iterations of the while loop. It thus remains to show that when this occurs, the following two conditions required by Lemma 3.11 are both satisfied:
@@
    Algorithm 2 terminates after $k\leq\kappa=O((\log L)/\varepsilon^{2})$ updates and returns a function $s:\{0,1\}^{n}\to[-1,1]^{L}$ that is simultaneously $(\mathcal{A},\varepsilon,\delta^{\prime})$-indistinguishable from $g^{*}$ and $(\mathcal{B},\varepsilon,\delta^{\prime})$-calibrated w.r.t. $g^{*}$.
@@
    Algorithm 2 runs in time $\mathsf{poly}(m,T,L,1/\varepsilon,\log(1/\delta))$, and for every $x\in\{0,1\}^{n}$, the function value $s(x)$ can be evaluated in time $\mathsf{poly}(m,T,L,1/\varepsilon,\log(1/\delta))$.
To prove that the output function $s=\hat{g}_{k}$ is $(\mathcal{A},\varepsilon,\delta^{\prime})$-indistinguishable from $g^{*}$ and $(\mathcal{B},\varepsilon,\delta^{\prime})$-calibrated as require by Item 1, it suffices to show that (35) and (36) are both violated. By (33) and (34), it suffices to show that the conditions in the “if” clauses at Algorithms 2 and 2 are both violated. This is indeed true for the final output $s=\hat{g}_{k}$ because otherwise Algorithm 2 would proceed to the next iteration of the “while” loop. Similarly, by (33), if the condition at Algorithm 2 is satisfied, we have
@@
By (34), if the condition at Algorithm 2 is satisfied, we have
@@
This allows us to prove that Algorithm 2 terminates after $k\leq\kappa=O((\log L)/\varepsilon^{2})$ updates following the same idea in the proof of Lemma 3.2, completing the proof of Item 1. Also, Item 2 can be proved similarly to Lemma 3.3 using the fact that both $\mathcal{A}$ and $\mathcal{B}$ run in time $T$ and their output functions $f$ and $r$ can be evaluated also in time $T$, according to Definitions 3.6 and 3.7. We omit the details. ∎

### v2

@@ 2.2 Mirror Descent
For $\lambda\geq 0$ and a norm $\|\cdot\|$ on $\mathbb{R}^{L}$, we say a convex function $\psi:\mathbb{R}^{L}\to\mathbb{R}$ is $\lambda$-smooth w.r.t. $\|\cdot\|$ if for every $h_{0},h\in\mathbb{R}^{L}$, letting $g_{0}$ be a subgradient of $\psi$ at $h_{0}$, we have
@@
Let $\varphi:\mathcal{G}\to\mathbb{R}$ be a closed convex function on a bounded convex set $\mathcal{G}\subseteq\mathbb{R}^{L}$, and let $\psi:\mathbb{R}^{L}\to\mathbb{R}$ be its convex conjugate. Assume that $\psi$ is $\lambda$-smooth w.r.t. norm $\|\cdot\|$ for some $\lambda\geq 0$. Assume conjugate pairs $(g_{k},h_{k}),(g_{k+1},h_{k+1})\in\mathcal{G}\times\mathbb{R}^{L}$ satisfy (23) for some $f\in\mathbb{R}^{L}$. Then for every $g^{*}\in\mathcal{G}$,
@@ 2.3 The Multiplicative Weights Algorithm
We give a proof of Lemma˜2.3 using Pinsker’s inequality in Section˜A.1.
@@ 3 Regularity Lemma with Weight-Restricted Calibration
As we discuss in Section˜1.5, the key to proving our main theorems (Theorems˜1.10 and 1.11) is to establish an enhanced regularity lemma that allows us to construct a low complexity function $s:\{0,1\}^{n}\to\Delta_{L}$ satisfying both multiaccuracy (12) and weight-restricted calibration (11). We formally state and prove this result in both the non-uniform and uniform settings.
@@ 3.1 Non-uniform Setting
We prove Theorem˜3.1 by constructing $s$ using Algorithm˜1 based on the Multiplicative Weights technique discussed in Section˜2.3. In particular, our analysis tracks the Fenchel-Young divergence $\Gamma=\Gamma_{\varphi,\psi}$ defined by the negative Shannon entropy $\varphi$ in (25) and its convex conjugate $\psi$ in (26). For every $g:\{0,1\}^{n}\to\Delta_{L}$ and $h:\{0,1\}^{n}\to\mathbb{R}$, we define
@@
For any norm $\|\cdot\|$ on $\mathbb{R}^{L}$ and any function $h:\{0,1\}^{n}\to\mathbb{R}^{L}$, we define
@@
As we will show, Theorem˜3.1 follows from Lemmas˜3.2 and 3.3 below.
@@ 18    $\mathtt{updated}\leftarrow\mathtt{TRUE}$;
Algorithm˜1 terminates after $k=O((\log L)/\varepsilon^{2})$ updates. Moreover, the output $s$ produced by the algorithm satisfies
@@
Therefore, (28) implies that $k=O((\log L)/\varepsilon^{2})$. Moreover, when Algorithm˜1 terminates, both conditions at Algorithms˜1 and 1 are violated, implying (27).
It remains to prove (28). We will prove the stronger claim that whenever we update from $h_{k}$ to $h_{k+1}$ at Algorithm˜1 or Algorithm˜1,
@@
We focus on the case where the update occurs at Algorithm˜1, and the other case of Algorithm˜1 can be handled similarly. By Lemma˜2.2, Lemma˜2.3 and the fact that $\|f\|_{\infty}\leq 1$, we have
@@
By the condition at Algorithm˜1 and the fact that $\|\hat{g}_{k}-g_{k}\|_{1}\leq\varepsilon/10$, we have
@@
There exists an implementation of Algorithm˜1 such that the following holds. Assume that the classes $\mathcal{F}$ and $\mathcal{R}$ both consist of functions with circuit complexity at most $T$. Then for every $k$, the circuit complexity of $\hat{g}_{k}$ is at most $(k+1)(T+L\cdot\mathsf{poly}(1/\varepsilon,\log L))$, and each coordinate of $\hat{g}_{k}(x)$ for every $x\in\{0,1\}^{n}$ is represented in binary format with $O(\log(L/\varepsilon))$ bits.
We will use the following lemma to prove Lemma˜3.3:
@@
[…]
Since each $p_{y}$ can be computed in circuit complexity $\mathsf{poly}(B,\log(1/\varepsilon))$, we have that $v$ can be computed in circuit complexity $L\cdot\mathsf{poly}(B,\log(1/\varepsilon),\log L)$. By (32) and the Lipschitzness of $\mathsf{softmax}$ (Lemma˜2.3), we have
@@
We are now ready to prove Lemma˜3.3.
Proof of Lemma˜3.3.
Let $\kappa=O((\log L)/\varepsilon^{2})$ be an upper bound on the total number of updates in Algorithm˜1, as guaranteed by Lemma˜3.2. Throughout the algorithm, we maintain a circuit $C_{k}$ that computes $x\mapsto(\hat{h}_{k}(x),\hat{g}_{k}(x))$ where $\hat{h}_{k}:\{0,1\}^{n}\to\mathbb{R}^{L},\hat{g}_{k}:\{0,1\}^{n}\to\Delta_{L}$ satisfy the following:
@@
    $\|\hat{g}_{k}-g_{k}\|_{\infty}\leq\varepsilon/10$ as required by Algorithms˜1, 1 and 1;
@@
We prove this by induction on $t$. When $t=0$, the algorithm initializes $h_{0}$ to be the constant zero function, so $\hat{h}_{0}$ can be computed in circuit complexity $O(L)$. By Lemma˜3.4, $\hat{g}_{0}$ can be computed in circuit complexity $L\cdot\mathsf{poly}(\log(1/\varepsilon),\log L)$.
Now suppose we have a circuit $C_{k}$ with size at most $T_{k}$ that computes $x\mapsto(\hat{h}_{k}(x),\hat{g}_{k}(x))$ satisfying the five conditions above, and let us construct the circuit $C_{k+1}$ that computes $x\mapsto(\hat{h}_{k+1}(x),\hat{g}_{k+1}(x))$. In Algorithm˜1, $h_{k+1}$ may be updated from $h_{k}$ at either Algorithm˜1 (using $f$) or Algorithm˜1 (using $r\circ\hat{g}_{t}$). Both cases can be handled similarly using our assumption that both $f$ and $r$ have circuit complexity at most $T$. We focus on the first case for simplicity. We first compute $\hat{f}$ with $\|\hat{f}\|_{\infty}\leq\varepsilon$ such that $\|\hat{f}-\varepsilon f\|_{\infty}\leq\varepsilon/30\kappa$, where each coordinate of $\hat{f}(x)$ need only have $O(\log(\kappa/\varepsilon))$ bits in binary representation. This can be done in circuit complexity $T+L\cdot\mathsf{poly}(\log(\kappa/\varepsilon))$, noting that we only need to read $O(\log(\kappa/\varepsilon))$ bits from each coordinate of $f(x)$. Then we subtract $\hat{f}$ from $\hat{h}_{k}$ to obtain $\hat{h}_{k+1}$. The subtraction takes circuit complexity $L\cdot\mathsf{poly}(\log(\kappa/\varepsilon))$ and ensures that $h_{k+1}$ satisfies conditions 1,2, and 4. Finally, by Lemma˜3.4, we can compute $\hat{g}_{k+1}$ satisfying conditions 3 and 5 from $\hat{f}_{k+1}$ in circuit complexity $O(L\cdot\mathsf{poly}(\kappa\varepsilon,\log(1/\varepsilon),\log L))$. The total circuit complexity of $C_{k+1}$ is
@@
We now complete the proof of Theorem˜3.1 using Lemma˜3.2 and Lemma˜3.3.
Proof of Theorem˜3.1.
Lemma˜3.2 ensures that the output function $s$ satisfies Conditions 2 and 3 in Theorem˜3.1 (indistinguishability and weight-restricted calibration). By Lemma˜3.3, we can make sure that $s$ has circuit complexity at most
@@
where we used the fact that $k+1=O((\log L)/\varepsilon^{2})$ by Lemma˜3.2. This establishes Condition 1 required by Theorem˜3.1. ∎
@@ 3.2 Uniform Setting
In Definitions˜3.8 and 3.9 above, the success probability of an oracle can be easily amplified by independent repetition. Suppose we have an oracle with success probability $\alpha>0$. We can construct an oracle with success probability at least $1-\delta$ by independently running the oracle $u=O(\alpha^{-1}\log(1/\delta))$ times to ensure that with probability at least $1-\delta/2$, at least one of the runs is successful. We can then perform a simple test using $O(\varepsilon^{-2}\log(u/\delta)))$ additional fresh data points to select a successful run with probability at least $1-\delta/2$ (with a constant-factor loss in $\varepsilon$). The overall success probability is at least $1-\delta$ by the union bound.
@@ 26    $\mathtt{updated}\leftarrow\mathtt{TRUE}$;
We prove Theorem˜3.10 using Algorithm˜2 as the simulating algorithm $\mathcal{S}$ based on the same multiplicative weights idea as Algorithm˜1. The difference is that instead of directly looking for $f$ and $r$ as in Lines 1 and 1 of Algorithm˜1, we obtain them efficiently by invoking the oracles $\mathcal{A}$ and $\mathcal{B}$ at Lines 2 and 2 in Algorithm˜2. Recall that for $y\in[L]$, we use $\mathbf{e}_{y}$ to denote the standard basis vector whose $y$-th coordinate is one. This ensures that for every $v\in\Delta_{L}$, we have $\mathbb{E}_{y\sim v}[\mathbf{e}_{y}]=v$.
Theorem˜3.10 follows immediately from the following lemma about Algorithm˜2:
@@
For every $\varepsilon\in(0,1/2)$, there exist $\kappa=O((\log L)/\varepsilon^{2}),\delta^{\prime}=\Omega(\delta/\kappa),m^{\prime}=O((\log(1/\delta^{\prime}))/\varepsilon^{2})$ such that when we run Algorithm˜2 with parameters $(\varepsilon,m^{\prime})$, with probability at least $1-\delta$, the following holds. Algorithm˜2 terminates after $k\leq\kappa$ updates and returns a function $s:\{0,1\}^{n}\to[-1,1]^{L}$ that is simultaneously $(\mathcal{A},\varepsilon,\delta^{\prime})$-indistinguishable from $g^{*}$ and $(\mathcal{B},\varepsilon,\delta^{\prime})$-calibrated w.r.t. $g^{*}$. Also, Algorithm˜2 runs in time $\mathsf{poly}(m,T,L,1/\varepsilon,\log(1/\delta))$, and for every $x\in\{0,1\}^{n}$, the function value $s(x)$ can be evaluated in time $\mathsf{poly}(m,T,L,1/\varepsilon,\log(1/\delta))$.
@@
By the Chernoff bound, choosing $m^{\prime}=O((\log(1/\delta^{\prime}))/\varepsilon^{2})$ is sufficient to ensure that each time Algorithm˜2 is performed, with probability at least $1-\delta^{\prime}$ we have
@@
Similarly, each time when Algorithm˜2 is performed, with probability at least $1-\delta^{\prime}$ we have
@@
By the definition of $(\mathcal{A},\varepsilon,\delta^{\prime})$-distinguishability (Definition˜1.6), each time when Algorithm˜2 is performed with $\hat{g}_{k}$ being $(\mathcal{A},\varepsilon,\delta^{\prime})$-distinguishable, with probability at least $1-\delta^{\prime}$ we have
@@
Similarly, by the definition of $(\mathcal{B},\varepsilon,\delta^{\prime})$-miscalibration Definition˜3.9, each time when Algorithm˜2 is performed with $\hat{g}_{k}$ being $(\mathcal{B},\varepsilon,\delta^{\prime})$-miscalibrated, with probability at least $1-\delta^{\prime}$ we have
@@
By the union bound, we can choose $\delta^{\prime}=\Omega(\delta/\kappa)$ to ensure that with probability at least $1-\delta$, all of the above are satisfied simultaneously in the first $\kappa+1$ iterations of the while loop. It thus remains to show that when this occurs, the following two conditions required by Lemma˜3.11 are both satisfied:
@@
    Algorithm˜2 terminates after $k\leq\kappa=O((\log L)/\varepsilon^{2})$ updates and returns a function $s:\{0,1\}^{n}\to[-1,1]^{L}$ that is simultaneously $(\mathcal{A},\varepsilon,\delta^{\prime})$-indistinguishable from $g^{*}$ and $(\mathcal{B},\varepsilon,\delta^{\prime})$-calibrated w.r.t. $g^{*}$.
@@
    Algorithm˜2 runs in time $\mathsf{poly}(m,T,L,1/\varepsilon,\log(1/\delta))$, and for every $x\in\{0,1\}^{n}$, the function value $s(x)$ can be evaluated in time $\mathsf{poly}(m,T,L,1/\varepsilon,\log(1/\delta))$.
To prove that the output function $s=\hat{g}_{k}$ is $(\mathcal{A},\varepsilon,\delta^{\prime})$-indistinguishable from $g^{*}$ and $(\mathcal{B},\varepsilon,\delta^{\prime})$-calibrated as require by Item˜1, it suffices to show that (35) and (36) are both violated. By (33) and (34), it suffices to show that the conditions in the “if” clauses at Algorithms˜2 and 2 are both violated. This is indeed true for the final output $s=\hat{g}_{k}$ because otherwise Algorithm˜2 would proceed to the next iteration of the “while” loop. Similarly, by (33), if the condition at Algorithm˜2 is satisfied, we have
@@
By (34), if the condition at Algorithm˜2 is satisfied, we have
@@
This allows us to prove that Algorithm˜2 terminates after $k\leq\kappa=O((\log L)/\varepsilon^{2})$ updates following the same idea in the proof of Lemma˜3.2, completing the proof of Item˜1. Also, Item˜2 can be proved similarly to Lemma˜3.3 using the fact that both $\mathcal{A}$ and $\mathcal{B}$ run in time $T$ and their output functions $f$ and $r$ can be evaluated also in time $T$, according to Definitions˜3.6 and 3.7. We omit the details. ∎